A source emits 31.4 W of radiant flux distributed uniformly in all directions. The luminous efficient is 60° lumen watt-1. What is the luminous intensity of the source?

Grade:upto college level

2 Answers

Kevin Nash
9 years ago
. Radiant flux = 31.4W, Solid angle = 4π Luminous efficiency = 60 lumen/W So, Luminous flux = 60 × 31.4 lumen And luminous intensity = Lu min ous Flux/4π = 60 * 31.4/4π = 150 candela
Deepak Patra
9 years ago
Sol. Radiant flux = 31.4W, Solid angle = 4π Luminous efficiency = 60 lumen/W So, Luminous flux = 60 × 31.4 lumen And luminous intensity = Lu min ous Flux/4π = 60 * 31.4/4π = 150 candela